An alloy of tin is 15% tin and weighs 20 pounds. A second alloy is 10% tin. How many pounds of the second alloy must be added to the first to get a 12% mixture?

Answer by mananth(16949)

------ percent ---------------- quantity
Alloy I 15 ----------------20 pounds
Alloy II 10 ----------------x pounds
Total 12----------------20+x pounds

15*20+10*x=12(20+x)
300+10x=240+12x
10x-12x=240-300
-2x=-60
/-2
x= 30 pounds 10 % Alloy II